6. (5 points) The management of Petro Garcia Inc. was discussing whether certain equipment should be written off as a charge6. (5 points) The management of Petro Garcia Inc. was discussing whether certain equipment should be written off as a charge to current operations because of obsolescence. This equipment has a cost of $900,000 with depreciation to date of $400,000 as of December 31, 2014. On December 31, 2014, management projected its future net cash flows from this equipment to be $300,000 and its fair value to be $270,000. The company intends to use this equipment in the future. Instructions (a) Prepare the journal entry (if any) to record the impairment at December 31, 2014.
Answer #1

(a) Carrying value = $900,000 - 400000 = $50000

Impairment loss = Carrying value - Recoverable value

= 500000 - 270000 = $230,000

Date Accounts Debit Credit
December 31, 2014 Loss on Impairment $230,000
Accumulated depreciation - equipment $230,000